# OATS: Outlier-Aware Pruning through Sparse and Low Rank Decomposition

### v1

@@ Abstract
The recent paradigm shift to large-scale foundation models has brought about a new era for deep learning that, while has found great success in practice, has also been plagued by prohibitively expensive costs in terms of high memory consumption and compute. To mitigate these issues, there has been a concerted effort in post-hoc neural network pruning techniques that do not require costly retraining. Despite the considerable progress being made, existing methods often exhibit a steady drop in model performance as the compression increases. In this paper, we present a novel approach to compressing large transformers, coined OATS, that utilizes the second moment information in the input embeddings to decompose the model weights into a sum of sparse and low-rank matrices. Without any retraining, OATS achieves state-of-the-art performance when compressing models by up to $60\%$ on large language models such as Llama-3 and Phi-3 and vision transformers such as ViT and DINOv2 while delivering up to 1.37×1.37$\times$1.37 × the CPU acceleration versus a model that was comparably pruned.
@@ 1 Introduction
Large scale transformer-based models have found great success in a number of domains ranging from image classification (Wu et al., 2020), language modelling (Devlin et al., 2019), and question answering (Brown et al., 2020). However, these models contain billions of parameters making them computationally expensive to train and deploy which has lead to an increased demand for resource-saving techniques like model quantization (Dettmers et al., 2022; Egiazarian et al., 2024), parameter efficient fine-tuning (Hu et al., 2022; Zhao et al., 2024b), and, most relevant to this work, neural network pruning (Frantar & Alistarh, 2023).
Pruning has been a key focus for model compression since the early days of deep networks (Mozer & Smolensky, 1988; LeCun et al., 1989; Hassibi & Stork, 1992). Over the years, various pruning techniques have emerged, introducing sparsity either before training (Lee et al., 2019; Wang et al., 2020; Tanaka et al., 2020; de Jorge et al., 2021), during training (Zhu & Gupta, 2018; Evci et al., 2020), or post-training (Benbaki et al., 2023). In the context of large foundation models, post-training pruning methods, particularly those requiring minimal (Xia et al., 2022; Ma et al., 2023) or no re-training (Frantar & Alistarh, 2023; Sun et al., 2024b; Ashkboos et al., 2024; Zhang et al., 2024b) are preferred for their computational efficiency. These techniques, when compressing models by $50\%$, have demonstrated the ability to accelerate end-to-end CPU inference by up to 1.8×1.8$\times$1.8 × (Frantar & Alistarh, 2023; Yin et al., 2024b) and GPU inference by up to 1.63×1.63$\times$1.63 × using structured N:M sparsity (Mishra et al., 2021), highlighting their potential in real-world applications.
Despite the significant advancements in pruning techniques, it was recently shown that current methods suffer from a consistent degradation in model performance as compression levels increase (Yin et al., 2024a). Moreover, although structured pruning offers greater potential for acceleration compared to unstructured pruning, it often imposes a much steeper trade-off in terms of model accuracy and effectiveness (Chen et al., 2022). These challenges underscore the need for more sophisticated pruning strategies to strike a better balance between efficiency and performance.
@@ 1.1 Contributions
To mitigate these issues, we introduce Outlier-Aware Pruning Through Sparse and Low Rank Decomposition (OATS), a novel retraining-free method for compressing large transformers that utilizes the second moment of the input embeddings to approximate the model’s weight matrices as a sum of a sparse matrix and a low-rank matrix.
We evaluate OATS on recent large language models – Phi-3 (Abdin et al., 2024) and Llama-3 (Dubey et al., 2024) – and vision transformers – ViT (Wu et al., 2020) and DinoV2 (Oquab et al., 2023) – demonstrating that OATS achieves new state-of-the-art performance across a wide range of commonly employed performance metrics. Furthermore, by combining structured pruning with unstructured pruning, our benchmarks utilizing the Deepsparse Engine by NeuralMagic (2021) reveal that OATS leads to more CPU speed-up across all levels of compression when compared to models that utilize just unstructured pruning.
To facilitate the interpretability of the sparse and low-rank terms found by OATS, we run experiments where we divide the compressed model into two models, a sparse model and a low-rank model, and visualize their respective attention heat maps utilizing attention rollout (Abnar & Zuidema, 2020). Our visualizations on vision transformers (Wu et al., 2020) reveal a complementary relationship between the two models, with each focusing on different key areas of the image, effectively segmenting it into distinct regions.
@@ 2 The OATS Algorithm
The key observation behind the OATS algorithm is that the weight matrices, ${\bm{W}}\in\mathbb{R}^{d_{out}\times d_{in}}$, in a transformer model can be faithfully approximated as a summation of a sparse and low rank matrix by solving the following optimization problem, commonly known as Robust PCA (Chandrasekaran et al., 2011; Candès et al., 2011):
@@ 2.1 Alternating Thresholding
To solve Equation 1, OATS leverages the alternating thresholding algorithms proposed by Netrapalli et al. (2014) and Bertsimas et al. (2024) that iteratively alternates between solving for the low-rank term ${\bm{L}}$, through singular-value thresholding, and for the sparse term ${\bm{S}}$, through hard-thresholding. Given a matrix ${\bm{A}}\in\mathbb{R}^{m\times n}$, singular-value thresholding, also known as truncated SVD, is defined as:
@@
where ${\bm{M}}\in\mathbb{R}^{m\times n}$ is a binary matrix with $k$ non-zero entries coinciding with the $k$ largest entries in magnitude in ${\bm{A}}$. Alternatively, the hard-thresholding can also be performed row-wise rather than layer-wise in which case ${\bm{M}}$ would be a binary matrix with $m\cdot\left\lfloor\frac{k}{m}\right\rfloor$ non-zero entries coinciding with the $\left\lfloor\frac{k}{m}\right\rfloor$ largest entries in magnitude in each row of ${\bm{A}}$.
These steps are summarized in Algorithm 1 on the right. To optimize memory usage, the low-rank term ${\bm{L}}$ is stored as its two low-rank components: ${\bm{U}}_{r}$ and $\mathbf{\Sigma}_{r}{\bm{V}}_{r}^{\top}$.
@@ 2.2 Incorporating Outlier Information
The alternating thresholding on its own yields suboptimal results because the activations of large-scale transformers exhibit a small number of large-magnitude features and altering these negatively impacts model performance (Kovaleva et al., 2021; Dettmers et al., 2022; Sun et al., 2024a). OATS takes inspiration from Wanda (Sun et al., 2024b) and incorporates outlier information through a diagonal scaling matrix ${\bm{D}}\in\mathbb{R}^{d_{in}\times d_{in}}$ that captures the second moment of the input activations
${\bm{D}}=\text{diag}\left(\sqrt{{\bm{X}}^{\top}{\bm{X}}}\right),$
where ${\bm{X}}\in\mathbb{R}^{B\times d_{in}}$ and $B$ is the product of the batch size and sequence length. The weight matrix ${\bm{W}}$ is then scaled by ${\bm{D}}$ before applying AlternatingThreshold:
${\bm{S}},{\bm{L}}=\textsc{AlternatingThreshold}\left({\bm{W}}{\bm{D}},N,r,k\right).$
To undo the scaling, OATS applies the inverse transformation to reach the compressed weight
@@
where it leverages the fact that ${\bm{D}}$ is diagonal so that it both preserves the sparsity pattern of ${\bm{S}}$ and is easy to invert. Aligned with (Frantar & Alistarh, 2023; Sun et al., 2024b), and Zhang et al. (2024b), the activations are calculated through a calibration set that is propagated through the compressed layers.
@@ 2.3 OATS Parameters
$\kappa=\frac{r(d_{out}+d_{in})}{(1-\rho)d_{out}\cdot d_{in}}.$
@@
2:      Layer Inputs: ${\bm{X}}^{\ell}\in\mathbb{R}^{B\times d_{in}}$
@@
9:  ${\bm{D}}\leftarrow\text{diag}(\sqrt{{\bm{X}}^{\top}{\bm{X}}})$
10:  ${\bm{L}},{\bm{S}}\leftarrow\textsc{AlternatingThreshold}({\bm{W}}{\bm{D}},N,r,k)$
@@ 3.1 Experiment Setup
We evaluate OATS on two state-of-the-art families of large language models: Phi-3 (Abdin et al., 2024) and Llama-3 (Dubey et al., 2024). To gauge the algorithm’s performance under various model sizes, we select Phi-3 Mini, a $3.8$B parameter model, Phi-3 Medium, a $14$B parameter model, and Llama-3 8B, an $8$B parameter model, so that models of various sizes are represented in our experiments. We utilize LM Harness developed by Gao et al. (2024) to evaluate zero-shot performance on eight tasks, five-shot performance on the Massive Multitask Language Understanding benchmark by Hendrycks et al. (2021), and language generation on WikiText-2.
@@
We benchmark our algorithm across a wide range of compression rates: {0.3, 0.4, 0.5, 0.6}. For compression rates at or below $0.5$, we compress all transformer blocks uniformly. At the higher compression rate of $0.6$, we utilize Outlier Weighed Layerwise Sparsity Ratios (OWL) developed by Yin et al. (2024b) which was shown to lead to significant performance improvements at higher compression rates. We exclude pruning any linear layers that are present in the model head and embeddings which conforms with what was done in the prior works by Frantar & Alistarh (2023), Sun et al. (2024b), and Zhang et al. (2024b).
@@
To measure the CPU speedup generated by the unstructured sparsity and low-rank compression of OATS, we utilize the DeepSparse Inference Engine developed by NeuralMagic (2021). More recently, NVIDIA’s sparse tensor cores are able to leverage for acceleration structured N:M sparsity, a sparsity pattern where every $N$ in $M$ parameters are zero (Mishra et al., 2021). We include in our experiments the natural extension of OATS to structured N:M sparsity where we impose the structured sparsity pattern onto the sparse term in the decomposition.
@@ 3.2 Results
Depicted in Table 2 below is the MMLU accuracy of OATS relative to current state of the art pruning algorithms. OATS is able to outperform all prior methods across all compression rates with an increasing gap as the compression rate increases. Specifically, at $50\%$ compression, OATS surpasses previous pruning algorithms by a margin of $2.52\%$ on Phi-3 Medium and $5.42\%$ on Phi-3 Mini.
Compression Rate Method Phi-3 Llama-3 8B
Mini Medium
0% Dense 70.34 76.78 64.97
30% SparseGPT 68.31 74.12 64.25
Wanda 67.63 75.18 63.67
DSNoT 68.02 75.13 63.72
OATS 68.84 76.15 65.22
40% SparseGPT 63.47 72.42 60.91
Wanda 64.15 73.34 60.33
DSNoT 63.57 73.20 59.99
OATS 65.75 74.99 62.46
50% SparseGPT 53.22 67.63 53.60
Wanda 54.57 69.76 49.83
DSNoT 54.28 68.65 49.20
OATS 59.99 72.28 56.46
@@
Table 3 below reports the zero-shot accuracy of OATS relative to current state of the art pruning algorithms averaged across the following eight popular tasks: PIQA (Bisk et al., 2020); HellaSwag (Zellers et al., 2019); Winogrande (Sakaguchi et al., 2021); OpenBookQA (Mihaylov et al., 2018); RTE (Wang et al., 2018); BoolQ (Clark et al., 2019); ARC-e and ARC-c (Clark et al., 2018). Mirroring the trend observed in the five-shot results, the improvement from OATS compared to prior pruning algorithms increases with compression, culminating in a $2.05\%$ advantage over prior methods at $50\%$ compression for Phi-3 Mini.
Compression Rate Method Phi-3 Llama-3 8B
Mini Medium
0% Dense 71.99 74.27 69.79
30% SparseGPT 70.63 74.53 69.08
Wanda 70.66 74.05 68.63
DSNoT 71.20 74.03 68.98
OATS 71.48 74.04 69.34
40% SparseGPT 69.18 74.40 67.58
Wanda 68.80 73.01 67.04
DSNoT 69.08 72.90 66.65
OATS 70.04 74.46 68.68
50% SparseGPT 66.36 73.25 64.66
Wanda 65.03 70.96 63.27
DSNoT 65.33 71.12 62.74
OATS 68.41 73.39 65.71
@@
Depicted in Table 4 below is the WikiText-2 perplexity of OATS relative to current state of the art pruning algorithms. At $50\%$ compression, OATS results in an $8.5\%$ reduction in perplexity on the larger Phi-3 Medium model, and an even larger $9\%$ reduction on Phi-3 Mini and Llama-3 8B.
Compression Rate Method Phi-3 Llama-3 8B
@@
0% Dense 9.50 6.21 10.17
30% SparseGPT 11.19 7.48 9.71
Wanda 10.71 7.28 9.39
DSNoT 10.51 7.11 9.36
OATS 10.27 6.85 9.59
40% SparseGPT 13.03 8.52 10.01
Wanda 12.59 8.49 9.74
DSNoT 12.17 8.24 9.60
OATS 11.53 7.70 9.24
50% SparseGPT 16.80 9.89 11.95
Wanda 17.23 10.12 12.36
DSNoT 16.68 9.96 12.41
OATS 15.18 9.05 10.87
Method Phi-3 Mini Llama-3 8B
SparseGPT 46.20 39.48
Wanda 44.22 31.20
DSNoT 44.75 33.28
OATS 52.59 44.46
@@
[…]
3.3 Ablations for OATS
We conduct ablation studies for OATS on Phi-3 Mini at $40\%$ compression rate with a rank ratio of $20\%$ to depict the quantitative impact of the following design choices that were made:

### v2

@@ Abstract
The recent paradigm shift to large-scale foundation models has brought about a new era for deep learning that, while has found great success in practice, has also been plagued by prohibitively expensive costs in terms of high memory consumption and compute. To mitigate these issues, there has been a concerted effort in post-hoc neural network pruning techniques that do not require costly retraining. Despite the considerable progress being made, existing methods often exhibit a steady drop in model performance as the compression increases. In this paper, we present a novel approach to compressing large transformers, coined OATS, that compresses the model weights by approximating each weight matrix as the sum of a sparse matrix and a low-rank matrix. Prior to the decomposition, the weights are first scaled by the second moment of their input embeddings, so as to ensure the preservation of outlier features recently observed in large transformer models. Without retraining, OATS achieves state-of-the-art performance when compressing large language models, such as Llama-3 and Phi-3, and vision transformers, such as Google’s ViT and DINOv2, by up to $60\%$, all while speeding up the model’s inference on a CPU by up to 1.37×1.37$\times$1.37 × compared to prior pruning methods. Our code is available at: https://github.com/stephenqz/OATS.
@@ 1 Introduction
Large scale transformer-based models have found great success in a number of domains ranging from image classification (Wu et al., 2020), language modeling (Devlin et al., 2019), and question answering (Brown et al., 2020). However, these models contain billions of parameters making them computationally expensive to train and deploy, which has lead to an increased demand for resource-saving techniques like model quantization (Dettmers et al., 2022; Egiazarian et al., 2024), parameter efficient fine-tuning (Hu et al., 2022; Zhao et al., 2024b), and, most relevant to this work, neural network pruning (Frantar & Alistarh, 2023).
Pruning has been a key focus for model compression since the early days of deep neural networks (Mozer & Smolensky, 1988; LeCun et al., 1989; Hassibi & Stork, 1992). Various pruning techniques have since emerged, introducing sparsity in the model parameters either before training (Lee et al., 2019; Wang et al., 2020; Tanaka et al., 2020; de Jorge et al., 2021), during training (Zhu & Gupta, 2018; Evci et al., 2020), or post-training (Benbaki et al., 2023). In the context of large foundation models, post-training pruning methods, particularly those requiring minimal (Xia et al., 2022; Ma et al., 2023) or no re-training (Frantar & Alistarh, 2023; Sun et al., 2024b; Ashkboos et al., 2024; Zhang et al., 2024b) are preferred for their computational efficiency. These techniques, when compressing models by $50\%$, have demonstrated the ability to accelerate end-to-end CPU inference by up to 1.8×1.8$\times$1.8 × (Yin et al., 2024b) and GPU inference by up to 1.63×1.63$\times$1.63 × using structured N:M sparsity (Mishra et al., 2021), highlighting their potential in reducing costs during deployment.
Despite the significant advancements in pruning techniques, it was recently shown that current methods suffer from a consistent degradation in model performance as compression levels increase (Yin et al., 2024a). Moreover, although structured pruning offers greater potential for acceleration compared to unstructured pruning, it often imposes a much steeper trade-off in terms of model accuracy and effectiveness (Chen et al., 2022). These challenges underscore the need for more sophisticated pruning strategies that can achieve better performance as compression increases.
@@ 1.1 Contributions
To mitigate these issues, we introduce Outlier-Aware Pruning Through Sparse and Low Rank Decomposition (OATS): a novel retraining-free method for compressing large transformers that approximates the model’s weight matrices as a sum of a sparse matrix and a low-rank matrix. In order to emphasize the outliers recently observed in large transformer models and preserve model performance (Kovaleva et al., 2021; Dettmers et al., 2022; Darcet et al., 2024; Sun et al., 2024a), OATS first scales the weights by the second moment of their corresponding input embeddings.
We evaluate OATS on recent large language models (LLMs) – Phi-3 (Abdin et al., 2024) and Llama-3 (Dubey et al., 2024) – and vision transformers – Google’s ViT (Wu et al., 2020) and DinoV2 (Oquab et al., 2023) – demonstrating that OATS achieves new state-of-the-art performance across a wide range of commonly employed performance metrics. Furthermore, by combining structured pruning with unstructured pruning, OATS accelerates CPU inference across all levels of compression when compared to models that utilize just unstructured pruning.
To gain a deeper understanding of the sparse and low-rank terms found by OATS, we split the compressed vision transformers (Wu et al., 2020) into two separate models, a sparse model and a low-rank model, and visualize their respective attention heat maps utilizing attention rollout (Abnar & Zuidema, 2020). These reveal a complementary relationship between the two models, with each focusing on different key areas of the image, effectively segmenting it into distinct regions.
@@ 2 The OATS Algorithm
The key observation behind the OATS algorithm is that the weight matrices, ${\bm{W}}\in\mathbb{R}^{d_{out}\times d_{in}}$, in a transformer model can be faithfully approximated as a summation of a sparse and low-rank matrix by solving the following optimization problem, commonly known as Robust PCA (Chandrasekaran et al., 2009; 2011; Candès et al., 2011):
@@ 2.1 Alternating Thresholding
To solve Equation 1, OATS leverages the alternating thresholding algorithms proposed by Zhou & Tao (2011), Netrapalli et al. (2014) and Bertsimas et al. (2024) that iteratively alternates between solving for the low-rank term ${\bm{L}}$, through singular-value thresholding, and for the sparse term ${\bm{S}}$, through hard-thresholding. Given a matrix ${\bm{A}}\in\mathbb{R}^{m\times n}$, singular-value thresholding, also known as truncated SVD, is defined as:
@@
where ${\bm{M}}\in\mathbb{R}^{m\times n}$ is a binary matrix with $k$ non-zero entries coinciding with the $k$ largest entries in magnitude in ${\bm{A}}$. These steps are summarized in Algorithm 1 on the right. To optimize memory usage, the low-rank term ${\bm{L}}$ is stored through its two low-rank components: ${\bm{U}}_{r}$ and $\mathbf{\Sigma}_{r}{\bm{V}}_{r}^{\top}$.
2.2 Alternative Sparsity Patterns
When performing the hard-threshold step, various restrictions can be enforced on the sparsity pattern of the sparse term of the decomposition for enhanced performance or speed-up. The following are two important cases:
Row-Wise Thresholding
The hard-thresholding can be performed row-wise rather than layer-wise in which case ${\bm{M}}$ would be a binary matrix with $m\cdot\left\lfloor\frac{k}{m}\right\rfloor$ non-zero entries coinciding with the $\left\lfloor\frac{k}{m}\right\rfloor$ largest entries in magnitude in each row of ${\bm{A}}$. Sun et al. (2024b) have shown this leads to better performance.
N:M Sparsity
The hard-thresholding can be applied at an even more granular level using N:M sparsity, where only the $N$ largest entries by magnitude in every group of $M$ entries in matrix ${\bm{A}}$ are nonzero. Recently, NVIDIA’s sparse tensor cores have been able to exploit such sparsity patterns for acceleration (Mishra et al., 2021).
@@ 2.3 Incorporating Outlier Information
The alternating thresholding on its own yields suboptimal results because the activations of large-scale transformers exhibit a small number of large-magnitude features and altering these (for example, through the sparse and low-rank approximation) negatively impacts model performance (Kovaleva et al., 2021; Dettmers et al., 2022; Darcet et al., 2024; Sun et al., 2024a). OATS takes inspiration from Wanda (Sun et al., 2024b) and computes a diagonal scaling matrix ${\bm{D}}\in\mathbb{R}^{d_{in}\times d_{in}}$ that captures the second moment of the input activations
${\bm{D}}=\sqrt{\text{diag}({\bm{X}}^{\top}{\bm{X}})},$
where ${\bm{X}}\in\mathbb{R}^{B\times d_{in}}$ and $B$ is the product of the batch size and sequence length. This diagonal matrix, containing large magnitudes for the outlier features, is used to amplify their significance in the reconstruction error of Equation 1, leading to the following alternative optimization problem:
$\min_{{\bm{S}},{\bm{L}}\in\mathbb{R}^{d_{out}\times d_{in}}}\|{\bm{W}}{\bm{D}}-{\bm{S}}-{\bm{L}}\|_{F}^{2}\;\text{ s.t. }\;\text{Rank}({\bm{L}})\leq r,\;\|{\bm{S}}\|_{0}\leq k.$
The solution of the problem is given by:
${\bm{S}},{\bm{L}}=\textsc{AlternatingThresholding}\left({\bm{W}}{\bm{D}},N,r,k\right)$
which gives a sparse plus low-rank approximation of ${\bm{W}}{\bm{D}}\approx{\bm{S}}+{\bm{L}}$. OATS then applies the inverse transformation to reach the final compressed weight:
@@
where it leverages the fact that ${\bm{D}}$ is diagonal so that it both preserves the sparsity pattern of ${\bm{S}}$ and is easy to invert. The original weight matrix is replaced with three matrices: the sparse matrix ${\bm{S}}{\bm{D}}^{-1}$, and two matrices coinciding with the low-rank factorization of ${\bm{L}}{\bm{D}}^{-1}$. Aligned with Frantar & Alistarh (2023); Sun et al. (2024b), and Zhang et al. (2024b), the activations are calculated through a calibration set that is propagated through the compressed layers.
@@ 2.4 OATS Parameters
$\kappa=\frac{\#\text{ of parameters in low-rank term}}{\#\text{ of nonzero parameters in compressed layer}}=\frac{r(d_{out}+d_{in})}{(1-\rho)d_{out}\cdot d_{in}}.$
@@
2:      Layer Inputs Propagated Through Prior Compressed Layers: ${\bm{X}}^{\ell}\in\mathbb{R}^{B\times d_{in}}$
@@
9:      ${\bm{D}}\leftarrow\sqrt{\text{diag}({\bm{X}}^{\top}{\bm{X}})}$
10:      ${\bm{L}},{\bm{S}}\leftarrow\textsc{AlternatingThresholding}({\bm{W}}{\bm{D}},N,r,k)$
@@ 3.1 Experiment Setup
We evaluate OATS on two state-of-the-art families of LLMs: Phi-3 (Abdin et al., 2024) and Llama-3 (Dubey et al., 2024). To gauge the algorithm’s performance under various model sizes, we select Phi-3 Mini, a $3.8$B parameter model, Phi-3 Medium, a $14$B parameter model, Llama-3 8B, an $8$B parameter model, and Llama-3 70B, a $70$B parameter model. We utilize LM Harness developed by Gao et al. (2024) to evaluate five-shot performance on the Massive Multitask Language Understanding benchmark by Hendrycks et al. (2021), zero-shot performance on eight tasks, and language generation on WikiText-2.
@@
We benchmark our algorithm across a wide range of compression rates: {0.3, 0.4, 0.5, 0.6}. For compression rates at or below $0.5$, we compress all transformer blocks uniformly. At the higher compression rate of $0.6$, we utilize Outlier Weighed Layerwise Sparsity Ratios (OWL) proposed by Yin et al. (2024b) which were shown to lead to significant performance improvements at higher compression rates. All linear layers in a transformer block are pruned uniformly to achieve the desired sparsity rate. We exclude pruning any linear layers that are present in the model head and embeddings which conforms with prior works by Frantar & Alistarh (2023), Sun et al. (2024b), and Zhang et al. (2024b).
@@
We benchmark the CPU speedup of OATS over its competitors using the DeepSparse Inference Engine developed by NeuralMagic (2021). For GPU speed-up, we include structured N:M sparsity experiments where the rank ratio is varied to measure the trade-off between compression and performance.
@@ 3.2 Results
Table 2, below, reports the MMLU accuracy of OATS relative to current state-of-the-art pruning algorithms. OATS is able to outperform all prior methods, across all compression rates, with an increasing gap as the compression rate increases. Notably, at $50\%$ compression, OATS surpasses previous pruning algorithms by a margin of $5.42\%$ on Phi-3 Mini, $2.52\%$ on Phi-3 Medium, $2.86\%$ on Llama-3 8B, and $2.03\%$ on Llama-3 70B.
Compression Method Phi-3 Llama-3
Mini ($\mathbf{3.8}$B) Medium ($\mathbf{14}$B) 8B 70B
0% Dense 70.34 76.78 64.97 79.63
30% SparseGPT 68.31 74.12 64.25 78.28
Wanda 67.63 75.18 63.67 79.15
DSNoT 68.02 75.13 63.72 79.00
OATS 68.84 76.15 65.22 78.47
40% SparseGPT 63.47 72.42 60.91 76.29
Wanda 64.15 73.34 60.33 77.16
DSNoT 63.57 73.20 59.99 77.70
OATS 65.75 74.99 62.46 77.89
50% SparseGPT 53.22 67.63 53.60 72.47
Wanda 54.57 69.76 49.83 72.04
DSNoT 54.28 68.65 49.20 72.76
OATS 59.99 72.28 56.46 74.79
@@
Table 3, below, reports the zero-shot accuracy of OATS relative to current state-of-the-art pruning algorithms averaged across the following eight commonly used tasks: PIQA (Bisk et al., 2020); HellaSwag (Zellers et al., 2019); Winogrande (Sakaguchi et al., 2021); OpenBookQA (Mihaylov et al., 2018); RTE (Wang et al., 2018); BoolQ (Clark et al., 2019); ARC-e and ARC-c (Clark et al., 2018). Mirroring the trend observed in the five-shot results, the improvement of OATS over prior pruning algorithms increases with compression, culminating in a $2.05\%$ advantage over prior methods when compressing Phi-3 Mini to $50\%$ of its size.
Compression Method Phi-3 Llama-3
Mini ($\mathbf{3.8}$B) Medium ($\mathbf{14}$B) 8B 70B
0% Dense 71.99 74.27 69.79 75.27
30% SparseGPT 70.63 74.53 69.08 75.07
Wanda 70.66 74.05 68.63 75.19
DSNoT 71.20 74.03 68.98 75.54
OATS 71.48 74.04 69.34 75.24
40% SparseGPT 69.18 74.40 67.58 74.63
Wanda 68.80 73.01 67.04 74.10
DSNoT 69.08 72.90 66.65 74.29
OATS 70.04 74.46 68.68 74.88
50% SparseGPT 66.36 73.25 64.66 73.17
Wanda 65.03 70.96 63.27 72.85
DSNoT 65.33 71.12 62.74 72.91
OATS 68.41 73.39 65.71 73.30
@@
Table 4, below, reports the WikiText-2 perplexity of OATS relative to current state-of-the-art pruning algorithms. At $50\%$ compression, OATS results in an $8.49\%$ reduction in perplexity on the larger Phi-3 Medium model, and an even larger $8.99\%$, $9.04\%$, and $9.30\%$ reduction on Phi-3 Mini, Llama-3 8B, and Llama-3 70B respectively.
Compression Method Phi-3 Llama-3
Mini ($\mathbf{3.8}$B) Medium ($\mathbf{14}$B) 8B 70B
0% Dense 9.50 6.21 10.17 2.68
30% SparseGPT 11.19 7.48 9.71 3.24
Wanda 10.71 7.28 9.39 3.28
DSNoT 10.51 7.11 9.36 3.27
OATS 10.27 6.85 9.59 3.07
40% SparseGPT 13.03 8.52 10.01 3.99
Wanda 12.59 8.49 9.74 4.08
DSNoT 12.17 8.24 9.60 4.10
OATS 11.53 7.70 9.24 3.68
50% SparseGPT 16.80 9.89 11.95 5.27
Wanda 17.23 10.12 12.36 5.38
DSNoT 16.68 9.96 12.41 5.58
OATS 15.18 9.05 10.87 4.78
Method Phi-3 Llama-3 8B
@@
SparseGPT 46.20 57.91 39.48
Wanda 44.22 58.49 31.20
DSNoT 44.75 58.20 33.28
OATS 52.59 64.10 44.46
@@
Table 5, on the right, is the 5-shot MMLU accuracy of models compressed to a higher compression rate of $60\%$, utilizing OWL ratios (Yin et al., 2024b). Following the trend above, OATS continues to outperforms prior methods by a margin of $6.39\%$ on Phi-3 Mini, $5.61\%$ on Phi-3 Medium, and $4.98\%$ on Llama-3 8B.
3.3 Studies and Hyperparameter Exploration
We conduct ablation studies for OATS, on Phi-3 Mini at $40\%$ compression rate with a rank ratio of $20\%$, to quantify the impact of the following design choices: